OpenStudy (anonymous):

HOW do u find the roots of x^2=8x

OpenStudy (anonymous):

x^2-8x=0 (x-4)^2-16=0

OpenStudy (anonymous):

Solutions x=0 x=8

OpenStudy (anonymous):

well that makes more sense :) thx